OpenStudy (anonymous):

Will give medal to the most helpful!!!!! Solve the system of equations. 2x+3y=1 y=3x+15

OpenStudy (anonymous):

I would use substitution so, 2x + 3(3x + 15) = 1

OpenStudy (anonymous):

you plug in what y equals for y in the other equation and solve for x, then plug that in and it will give you y

OpenStudy (anonymous):

arrange the equations by putting it equal to zero cencel out x or y to get the value of other then put that value in any eq to get the remaining value

OpenStudy (anonymous):

I can spell it out step by step if you want mad

OpenStudy (anonymous):

ok......I know the answer should be an ordered pair because the solution to the equations is where the two lines meet

OpenStudy (anonymous):

so 1 would be the y value?

OpenStudy (anonymous):

2x + 3y = 1 y = 3x + 15 so we take the first equation and replace y with what we know it is equal to from 2nd eq - 2x + 3(3x + 15) = 1 we solve for x - x = -4 we plug that into an original eq for example the second one- y = 3(-4) + 15 so, y = 3 our ordered pair is (-4,3)

OpenStudy (anonymous):

ohhhhhh ok

OpenStudy (anonymous):

we can check by plugging these into an equation y = 3x + 15 3 = 3(-4) + 15 3=3 True! 2x + 3y = 1 2(-4) + 3(3) = 1 1=1 True! So our ordered pair is correct

OpenStudy (anonymous):

thank you both SO much!

OpenStudy (anonymous):

no worries I hope that makes sense that is the substitution method you can also use elimination